Project

General

Profile

NDOS Commissioning Meeting 2011/09/01

Attendance: Jon, Mat, Chad, Mark Mathis, Enrique, Athans, Philip, Peter, Minerba, Jarek, Chris, Zeb, Nick, Jan, Susan, Leon, Gavin, Martin

Agenda:

  1. DAQ Update
  2. DCS Update
  3. Beam & Detector Update
  4. DataCheck Update
  5. BNB Event Update
  6. DSO on multiple channels simultaneously
  7. AOB

DAQ Update

  • Peter: nothing major; there were a few updates that the DAQ group wanted to implement but haven't had time. Perhaps tomorrow (before noon).
  • Jon: no timing issues? Peter: everything looks fine lately.

DCS Update

  • Martin:
    • smoke alarm was triggered during power outage, needs to be added to a checklist for when power outage occurs.
    • scripts will now check to make sure crate is powered on. Also, scripts have been cleaned up.

Beam & Detector Update

  • Jon: spoke to Chuck Andrews; he will send Jon a brief summary of what was done on Tues. morning, but basically said that in general the grounding looked ok, although there is room for improvement.
  • Mat: went to NDSB and tried to recreate DAQ fallover by touching detector and shuffling things around. Tried removing e-box cover, no problem. Tried running shop-vac, no problem. Ran shop-vac + removed e-box cover, and the DAQ fell over. Peter restarted the DAQ but they were unable to reproduce the effect.
  • Jon will let Chuck Andrews know about these tests.
  • Power will be shut off again on Tues. morning
  • Leon: All 6 of full APD assemblies sent back had leaks, and they were able to identify where they are. Got some heatsinks from IU, and 90% of those were found to have leaks. Found leak location in 28/30 heatsinks. 99% of leaks come from TECC lead holes and wires. Will try to pot these.
  • Jon: requested Leon provide a list of
  • Martin: FEB card that was removed (with bad FEB temp. readout) is in the mail to Harvard

Booster Timing Update

  • possibly sees a peak in last week's data.

DataCheck Update

  • NuMI peak looks good
  • Still lots of empty events in some runs
  • nearline is processing new subruns every ~hour. Webpage is updated every 5 minutes.

DSO on multiple channels simultaneously

AOB